OpenStudy (anonymous):

2/3a=-24 I don't want the answer.

OpenStudy (anonymous):

so what is your quenstion?

OpenStudy (freckles):

\[\frac{a}{b}x=\frac{c}{d} \\ \text{ so solve for x we need to multiply both sides by } \frac{b}{a} \\ \text{ doing this gives us: } \frac{b}{a} \cdot \frac{a}{b} x=\frac{b}{a} \cdot \frac{c}{d} \\ \text{ we see this cancels the } a's \text{ and the } b's \text{ on the left hand side } \\ x=\frac{bc}{ad}\]

OpenStudy (anonymous):

How do I solve it.

OpenStudy (freckles):

so what will you multiply both sides of your equation by?

OpenStudy (anonymous):

\[so jjdsjdjsjddjsjd\]

OpenStudy (xapproachesinfinity):

that is still mysterious the way you wrote it are we talking about \(\huge \frac{2}{3a}=-24\) or \(\huge \frac{2}{3}a=-24\)

OpenStudy (anonymous):

the second one

OpenStudy (xapproachesinfinity):

hmm then use what @freckles said